# Complements of discriminants of real parabolic function singularities. II

### v2

@@ Abstract.
We provide a complete list of the connected components of the spaces of non-discriminant functions within standard versal deformations of function singularities of classes $X_{9}$, $J_{10}$ and $P_{8}^{1}$ (as well as a partial list for the remaining class, $P_{8}^{2}$). Thus, we prove (and improve in one particular case) the corresponding conjectures from the previous work [18] with the same title. As an application, we enumerate all local Petrovskii lacunas near arbitrary parabolic singularities of wavefronts of hyperbolic PDEs. In particular, we discover a new local lacuna at the $P_{8}^{2}$ singularities. We also show that the complements of the discriminant varieties of $X_{9}^{+}$ and $P_{8}^{1}$ singularities have nontrivial one-dimensional homology groups, unlike all simple singularities.
@@ 1. Introduction
Discriminant varieties of smooth functions depending on parameters are the sets of parameters for which the zero-level sets of these functions are singular. These varieties appear in many problems of PDE theory and integral geometry, often under the name “wavefronts”, see e.g. [1], [2], [3], [8], [15][16], [17].
Many important functions in these theories arise as integral transforms. These functions are defined on the parameter spaces of certain function deformations (called generating families) and are equal to the integrals along the cycles in the zero-level sets of the corresponding functions. These integrals are regular outside the discriminant varieties of the deformations. Their analytic behavior near the discriminants depends on the singularity type of the deformed discriminant function and on the connected component of the complement of the discriminant. Thus, enumerating these components is one of the first steps in studying the analytical properties of these integral transforms.
@@
A function singularity is a germ of a regular analytic function $({\mathbb{R}}^{n},0)\to({\mathbb{R}},0)$ at its critical point. Two function singularities at points $a$ and $b$ are equivalent if they can be transformed to each other via a local diffeomorphism $({\mathbb{R}}^{n},a)\to({\mathbb{R}}^{n},b)$. Two function singularities, $f(x_{1},\dots,x_{n})$ and $\tilde{f}(\tilde{x}_{1},\dots,\tilde{x}_{m})$, are stably equivalent if they become equivalent after the summation with non-degenerate quadratic forms in additional variables, that is,
@@
An $l$-parametric deformation of a function singularity $f$ is a family of analytic functions, $f_{\lambda}$, where $\lambda$ runs through a neighborhood of the origin in ${\mathbb{R}}^{l}$, such that $f_{0}\equiv f$ and the function $F(x,\lambda)\equiv f_{\lambda}(x)$ in $n+l$ variables is regular analytic. The discriminant of the deformation $F$ is the set of parameter values $\lambda\in{\mathbb{R}}^{l}$ such that the corresponding function $f_{\lambda}$ has a real critical point with zero critical value (or, equivalently, the zero-level set $f_{\lambda}^{-1}(0)$ is singular near real space).
For each isolated function singularity, studying all its deformations reduces to studying any its sufficiently large (so-called versal, see [2]) deformation. The discriminant sets of all versal deformations of stably equivalent singularities are topologically equivalent, and the components of their complements are in a one-to-one correspondence. Therefore, for any stable equivalence class of singularities it is sufficient to enumerate the components for only one versal deformation of one function of this class.
The list of equivalence classes of function singularities begins with simple singularities, which were distinguished by V. Arnold, see [2]. All of these singularities are stably equivalent to functions in two variables, and the entire geometric study of their deformations and discriminants reduces to the study of functions in two variables. The components of discriminant complements for these singularities were listed earlier, see § 1.3.
Table 1. Normal forms of simple, $X_{9}$ and $J_{10}$ singularities in two variables and $P_{8}$ singularities in three variables
@@
In this work we list analogous components for the next natural class in the hierarchy of function singularities, the parabolic singularities, which were also distinguished by V. Arnold. The classification of complex parabolic singularities up to the stable equivalence consists of three classes $P_{8}$, $X_{9}$, and $J_{10}$. The classification of their real forms includes more classes: $P_{8}^{1}$, $P_{8}^{2}$, $X_{9}^{+}$, $X_{9}^{-}$, $X_{9}^{1}$, $X_{9}^{2}$, $J_{10}^{1}$, and $J_{10}^{3}$. Functions of classes $X_{9}$ and $J_{10}$ are stably equivalent to the functions in two variables, while the $P_{8}$ singularities are stably equivalent to the non-degenerate cubic forms in three variables. Any function singularity in two variables of class $X_{9}^{*}$ or $J_{10}^{*}$, as well as any function in three variables of one of classes $P_{8}^{*}$ can be reduced to a normal form indicated in Table 1 by an appropriate choice of local coordinates centered at the critical point. Additionally, this table provides the normal forms of simple singularities $A_{k},D_{k},$ and $E_{k}$, into which the parabolic singularities can split after small perturbations.
We give complete lists of components for all parabolic singularity classes except $P_{8}^{2}$, for which we have only partial results.
@@
O. Viro [24] found all topological types of non-discriminant perturbations of the $X_{9}$ and $J_{10}$ singularities. By definition, two small non-discriminant perturbations $f_{\lambda}:{\mathbb{R}}^{2}\to{\mathbb{R}}$ of the same isolated function singularity have the same topological type if there is an isotopy between their zero-level sets in a neighborhood of the critical point that preserves the behavior of these sets at the boundary of the neighborhood. The topological type is an invariant of the components of the discriminant complements. The “rigid isotopy” classification of parabolic singularity perturbations given in the present work is only slightly stricter than the topological classification, see Table 2. Counting the components of the discriminant complements of $P_{8}$ singularities is related to (though is different from) the isotopy classification of nonsingular cubic surfaces in ${\mathbb{R}}^{3}$, which is described in detail in [7]. This classification consists of five and ten classes for $P_{8}^{1}$ and $P_{8}^{2}$ singularities, respectively.
@@ 1.2. On the method
In [18], a system of invariants of connected components of the complements of the discriminant varieties of function singularities is described. Complete lists of possible values of these invariants for parabolic singularities are given in [18], and all of these values are realized by concrete polynomials. In the same paper, we also conjectured that the resulting lists of connected components are complete. These lists consist of 7, 15, 7, 14, 52, 11, and 33 components for the singularities of classes $P_{8}^{1}$, $P_{8}^{2}$, $X_{9}^{\pm}$, $X_{9}^{1}$, $X_{9}^{2}$, $J_{10}^{1}$, and $J_{10}^{3}$ respectively. Below, we prove that almost all of these lists are indeed complete. However, some two sets of perturbations of the $J_{10}^{1}$ singularities indicated in [18] consist of pairs of different connected components. Also, we present below partial results for the case $P_{8}^{2}$, which prove that at least 10 out of 15 classes listed in [18] are indeed represented by only one connected component each. Thus, the number 11 above is replaced in Table 2 by 13, and the number 15 is still under suspicion. These results are summarized in Theorems 5, 13, 18, 28, 31, 37, and 41 below.
@@ 1.3. On the discriminant complements for simple singularities
E. Looijenga [13] proved that all local components of discriminant complements of real simple singularities are topologically trivial. Conversely, we demonstrate that at least three such components of parabolic singularities of each of the classes $X_{9}^{+}$ and $X_{9}^{-}$, as well as one component of the $P_{8}^{1}$ class, have nontrivial one-dimensional homology groups; see Theorems 12 and 39.
@@
In the theory of hyperbolic PDEs, local Petrovskii lacunas are the local connected components of wavefront complements where wave functions behave regularly (see, for example, [1], [8], [16]). Near the non-singular points of wavefronts (corresponding to the Morse functions) the lacunas were enumerated in [6] and [5]. Near the simplest non-Morse points of classes $A_{2}$ and $A_{3}$, this was done in [8], and near all simple singularities in [15]. In [17], a list of Petrovskii lacunas at parabolic singular points of wavefronts was presented, and it was conjectured that the list is complete. We prove this conjecture for almost all parabolic classes but discover one additional local lacuna in the $P_{8}^{2}$ case, see § 5. For the singularities of classes $P_{8}^{1}$, $X_{9}^{*}$ and $J_{10}^{*}$, these results follow directly from our complete lists of components of discriminant complements. Our partial results on enumerating components for the $P_{8}^{2}$ singularities are sufficient to prove that our new list of local lacunas is also complete in that case.
@@ 2. Results for the $X_{9}^{*}$ classes
The standard versal deformations of functions in all these classes have the same form: they are families of polynomials
@@
depending on nine parameters $\lambda_{1},\dots,\lambda_{8},$ and $A$, where $f_{A}$ is the corresponding normal form from Table 1. According to versality theory (see, for example, [2]), studying all other deformations of singularities of stable equivalence classes $X_{9}^{*}$ can be reduced to studying these families.
@@
The parameters $\{\lambda_{1},\dots,\lambda_{8},A\}$ identify the spaces $\Phi_{+}$ and $\Phi_{1}$ with ${\mathbb{R}}^{8}\times(-1,1)$, and the space $\Phi_{2}$ with ${\mathbb{R}}^{8}\times(-\infty,-1)$. The discriminant subvarieties of these spaces consist of functions that have at least one real critical point with zero critical value. These varieties divide the spaces $\Phi_{*}$ into connected components. Enumerating of these components is our goal in this section.
@@ 2.2. Virtual functions and virtual components
The virtual function associated with a generic real polynomial is a collection of topological characteristics of the complex zero-level set of this polynomial, depending also on the collections of paths in ${\mathbb{C}}^{1}$ defining the bases of vanishing cycles in the homology of this level set. Specifically, the data of a generic polynomial $f(x,y)$ of some class $X_{9}^{*}$ consists of
@@
The (abstract) virtual functions of class $\Phi_{*}$ include all virtual functions associated with generic real polynomials of this class, as well as all analogous collections of data that can be obtained from them via standard flips that model surgeries of real functions. The formal graph of any type $X_{9}^{*}$ is the graph, whose vertices correspond to all the virtual functions of the corresponding class $\Phi_{*}$, and whose edges correspond to the standard flips. The virtual components of the formal graph are the connected subgraphs, into which the formal graph splits when all edges modeling the surgeries that cross the discriminant are removed. The set-valued invariant of a generic polynomial $f$ of class $\Phi_{*}$ is the virtual component of the formal graph containing the virtual functions associated with this polynomial. This virtual component is denoted by $S(f)$.
@@
Enumerating the virtual components of such a graph is a combinatorial problem that can be solved using a computer program. The work [18] provides the results of these computations for all parabolic singularities, as well as the immediate consequences of these computations in the terms of real functions. The remaining problem is to find all real connected components with this value of the set-valued invariant for each virtual component.
@@
b)) every virtual function from the virtual component $S(f)$ is associated with some generic real polynomial $\tilde{f}$ from the same connected component of the complement of the discriminant as $f$. $\Box$
@@
In the statement of Proposition 3 of [21], the functions $f$ and $\tilde{f}$ belong to the broader class of all degree-four Morse polynomials with nondegenerate principal homogeneous parts. However, the entire construction proving this proposition in [21] can be performed within the spaces of the versal deformations (1). Thus, it also proves the above Proposition 2.
These constructions enable us to estimate from above the number of isotopy classes of generic Morse perturbations associated with any given virtual function. These estimates are given in Propositions 8, 15, and 20 for classes $X_{9}^{+}$, $X_{9}^{1}$ and $X_{9}^{2}$ separately. Consequently, we obtain estimates of the number of connected components of the complement of a real discriminant which are associated with each virtual component. Specifically, in the cases $X_{9}^{+}$ and $X_{9}^{2}$, all real connected components of discriminant complements associated with the same virtual component can be obtained from each other by the action of a transformation group of order eight. In the case $X_{9}^{1}$ the analogous group is of order four. Then, for each virtual component in the list from [18], we compare all real components associated with this virtual component and determine which of these components are identical. This comparison provides the results of the present work concerning the $X_{9}$ singularities. The strategy for the $J_{10}$ and $P_{8}$ singularities is exactly the same.
@@
In [21], [22], and [23], we use the same strategy with a different definition of virtual components. In these works, we classify the Morse perturbations of function singularities up to isotopy within the space of Morse functions instead of the space of non-discriminant functions. Thus, the virtual components of the formal graph are determined there by removing the edges that change the Morse type.
@@
For any Morse polynomial $f:{\mathbb{R}}^{n}\to{\mathbb{R}}$, the number $\mbox{Ind}(f)$ equals the number of its real critical points with a negative critical value and an even Morse index minus the number of critical points with a negative critical value and with an odd Morse index.
It is easy to see that this number is an invariant of the connected components of the discriminant complements (and of associated virtual components).

### v3

@@ Abstract.
We list all connected components of sets of non-discriminant functions within versal deformations of all parabolic function singularities. These singularities are the second most common family of singularity classes of smooth functions after simple singularities. Thus, we prove (and improve in one particular case) the corresponding conjectures from the previous work [24] with the same title. As an application, we enumerate all local Petrovskii lacunas near arbitrary parabolic singularities of wavefronts of hyperbolic PDEs. We also show that the complements of the discriminant varieties of $X_{9}^{+}$ and $P_{8}^{1}$ singularities have nontrivial one-dimensional homology groups, in contrast to all simple singularities.
These results are applications of a general method for investigating non-singular perturbations of real function singularities. An important part of this method is a computer program that formalizes Picard–Lefschetz theory and surgeries of Morse functions.
@@ 1. Introduction
Discriminant varieties of families of smooth functions depending on parameters are the sets of parameters for which the zero-level sets of the corresponding functions are singular. These varieties appear in many problems of PDE theory and integral geometry, often under the name “wavefronts”, see e.g. [2], [3], [4], [9], [14], [20], [22], [23].
Many important functions in these theories arise as integral transforms. Their arguments are the parameters of certain function families (called generating families), and their values are defined by the integrals along the cycles in the zero-level sets of the corresponding functions. These integral functions are regular outside the discriminant varieties of the families. The analytic behavior of such a function near a discriminant point depends on the singularity type of the corresponding discriminant function and on the connected component of the complement of the discriminant, from which we approach this point. Thus, enumerating these components is one of the first steps in studying the analytical properties of the integral transforms.
@@
A function singularity is a germ of a regular analytic function $({\mathbb{R}}^{n},a)\to({\mathbb{R}},0)$ at its critical point. Two function singularities at points $a$ and $b$ are equivalent if they can be transformed to each other via a local diffeomorphism $({\mathbb{R}}^{n},a)\to({\mathbb{R}}^{n},b)$ (i.e., they have the same expression in appropriate local coordinate systems centered at $a$ and $b$). Two function singularities, $f(x_{1},\dots,x_{n})$ and $\tilde{f}(\tilde{x}_{1},\dots,\tilde{x}_{m})$, are stably equivalent if they become equivalent after the summation with non-degenerate quadratic forms in additional variables, that is,
@@
An $l$-parametric deformation of a function singularity $f$ is a family of analytic functions, $f_{\lambda}$, where $\lambda$ runs through a neighborhood of the origin point $0\in{\mathbb{R}}^{l}$, such that $f_{0}\equiv f$ and the function $F(x,\lambda)\equiv f_{\lambda}(x)$ in $n+l$ variables is regular analytic. The discriminant variety of the deformation $F$ is the set of parameter values $\lambda\in{\mathbb{R}}^{l}$ such that the corresponding function $f_{\lambda}$ has a real critical point with zero critical value near the origin in ${\mathbb{R}}^{n}$ (or, equivalently, the zero-level set of the analytic continuation of $f_{\lambda}$ to ${\mathbb{C}}^{n}$ is singular at this point).
For each isolated function singularity, studying all its deformations reduces to studying any its sufficiently large (so-called versal, see [3]) deformation. The discriminant varieties of all versal deformations of stably equivalent singularities are topologically equivalent, and the connected components of their complements in the parameter spaces are in a one-to-one correspondence. Therefore, for any stable equivalence class of singularities it is sufficient to enumerate these components for only one versal deformation of one function of this class.
The list of equivalence classes of function singularities begins with simple singularities, which were distinguished by V. Arnold, see [3]. All of these singularities are stably equivalent to functions in two variables, and the entire geometric study of their deformations and discriminants reduces to the study of such functions. The components of discriminant complements for simple singularities were listed earlier, see § 1.3.
Table 1. Normal forms of simple, $X_{9}$ and $J_{10}$ singularities in two variables and of $P_{8}$ singularities in three variables
@@
In this work we list all analogous components for the next natural collection of equivalence classes in the hierarchy of function singularities, the parabolic singularities, which were also distinguished by V. Arnold. The list of complex parabolic singularities up to the stable equivalence consists of three classes $P_{8}$, $X_{9}$, and $J_{10}$. The classification of their real forms includes more classes: $P_{8}^{1}$, $P_{8}^{2}$, $X_{9}^{+}$, $X_{9}^{-}$, $X_{9}^{1}$, $X_{9}^{2}$, $J_{10}^{1}$, and $J_{10}^{3}$. Singularities of classes $X_{9}$ and $J_{10}$ are stably equivalent to the functions in two variables, while the $P_{8}$ singularities are stably equivalent to the non-degenerate cubic forms in three variables. Any real function singularity in two variables of class $X_{9}^{*}$ or $J_{10}^{*}$, as well as any singularity in three variables of one of classes $P_{8}^{*}$, has the normal form indicated in Table 1 in appropriate local coordinates centered at the critical point. Additionally, this table provides the normal forms of simple singularities $A_{k},D_{k},$ and $E_{k}$, into which the parabolic singularities can split after small perturbations.
@@
O. Viro [30] found all topological types of non-discriminant perturbations of the $X_{9}$ and $J_{10}$ singularities. By definition, two small non-discriminant perturbations $f_{\lambda}:{\mathbb{R}}^{2}\to{\mathbb{R}}$ of the same isolated function singularity have the same topological type if there is an isotopy between their zero-level sets in a neighborhood of the critical point that preserves the behavior of these sets at the boundary of the neighborhood. Clearly, the topological type is an invariant of the connected components of the discriminant complements. Our “rigid isotopy” classification of non-discriminant perturbations separating all these components is only slightly stricter than the topological classification, see Table 2. Counting the components of the discriminant complements of $P_{8}$ singularities is related to (though is different from) the isotopy classification of nonsingular cubic surfaces in ${\mathbb{R}}^{3}$, which is described in detail in [8]. This classification consists of five and ten classes for $P_{8}^{1}$ and $P_{8}^{2}$ singularities, respectively.
@@ 1.2. On the method
In [24], a system of invariants of connected components of the complements of the discriminant varieties of function singularities is described. Complete lists of possible values of these invariants for parabolic singularities are given in [24], and all of these values are realized by concrete polynomials. In the same paper, we also conjectured that the resulting lists of connected components are complete. These lists consist of 7, 15, 7, 14, 52, 11, and 33 components for the singularities of classes $P_{8}^{1}$, $P_{8}^{2}$, $X_{9}^{\pm}$, $X_{9}^{1}$, $X_{9}^{2}$, $J_{10}^{1}$, and $J_{10}^{3}$ respectively. Below, we prove that almost all of these lists are indeed complete. However, some two sets of perturbations of the $J_{10}^{1}$ singularities indicated in [24] consist of pairs of different connected components. Thus, the number 11 above is replaced in Table 2 by 13. These results are summarized in Theorems 6, 14, 19, 29, 32, 38, and 42 below.
@@ 1.3. On the discriminant complements for simple singularities
E. Looijenga [16] proved that all local components of discriminant complements of real simple singularities are homotopically trivial. Conversely, we demonstrate that at least three such components of parabolic singularities of each of the classes $X_{9}^{+}$ and $X_{9}^{-}$, as well as one component of the $P_{8}^{1}$ class, have nontrivial one-dimensional homology groups; see Theorems 13 and 40.
@@
In the theory of hyperbolic PDEs, local Petrovskii lacunas are the local connected components of wavefront complements where wave functions behave regularly (see, for example, [2], [9], [14], [17], [22]). Near the non-singular points of wavefronts (corresponding to the Morse singularities of generating functions) the lacunas were enumerated in [7] and [6]. Near the simplest non-Morse points of classes $A_{2}$ and $A_{3}$, this was done in [9], and near all simple singularities in [20]. In [23], a list of Petrovskii lacunas at parabolic singular points of wavefronts was presented, and it was conjectured that the list is complete. We prove this conjecture for almost all parabolic classes but discover one additional local lacuna in the $P_{8}^{2}$ case, see § 5.
@@ 2. Results for the $X_{9}^{*}$ classes
The standard versal deformations of these singularities in all four cases have the same form: they are families of polynomials
@@
depending on nine parameters $\lambda_{1},\dots,\lambda_{8},$ and $A$, where $f_{A}$ is the corresponding normal form from Table 1. According to versality theory (see [3], § 8), studying all other deformations of singularities of stable equivalence classes $X_{9}^{*}$ can be reduced to studying these families. The entire study of $X_{9}^{-}$ singularities and their deformations can be reduced to that of $X_{9}^{+}$ via the multiplication by 1-1.
@@
The parameters $\{\lambda_{1},\dots,\lambda_{8},A\}$ identify the spaces $\Phi_{+}$ and $\Phi_{1}$ with ${\mathbb{R}}^{8}\times(-1,1)$, and the space $\Phi_{2}$ with ${\mathbb{R}}^{8}\times(-\infty,-1)$. The discriminant subvarieties of these spaces consist of functions that have at least one real critical point with zero critical value. These varieties divide the spaces $\Phi_{*}$ into connected components. Enumerating these components is our goal in this section.
@@ 2.2. Virtual functions and virtual components
A virtual function associated with a generic real polynomial $f$ is a collection of topological characteristics of the complex zero-level set of this polynomial. Specifically, the data of a generic polynomial $f(x,y)$ of some class $X_{9}^{*}$ consists of
  1. (1)
@@
Remark 1.
Besides the function $f$, a virtual function depends on the collection of paths in ${\mathbb{C}}^{1}$ defining the base of vanishing cycles in the homology of the level set. If the function has no more than one pair of non-real critical values, then there is only one canonical system of paths, and hence only one associated virtual function.
If all critical points of a polynomial are real, then the intersection indices from item (2) are easily determined by the remaining elements (1), (3), and (4) of a virtual function and hence may not be included in its description, see Theorem V.1.3 in [22].
The (abstract) virtual functions of class $\Phi_{*}$ include all virtual functions associated with generic real polynomials of this class, as well as all analogous collections of data that can be obtained from them via standard flips that model elementary surgeries of real functions. The formal graph of any type $X_{9}^{*}$ is the graph, whose vertices correspond to all the virtual functions of the corresponding class $\Phi_{*}$, and whose edges correspond to the standard flips. The virtual components of the formal graph are the connected subgraphs, into which the formal graph splits when all edges modeling the surgeries that cross the discriminant are removed. The set-valued invariant of a generic polynomial $f$ of class $\Phi_{*}$ is the virtual component of the formal graph containing the virtual functions associated with this polynomial. This virtual component is denoted by $S(f)$.
@@
Enumerating the virtual components of such a graph is a combinatorial problem. An (available) computer program solving it is described in [22] and [24]. The work [24] provides the results of these computations for all parabolic singularities, as well as the immediate consequences of these computations in the terms of real functions. The remaining problem is to find all real connected components associated with each virtual component.
@@
b)) for any generic polynomial $f$ of this class, every virtual function from the virtual component $S(f)$ is associated with some generic real polynomial $\tilde{f}$ from the same connected component of the complement of the discriminant as $f$. $\Box$
@@
In the statement of Proposition 3 of [27], the functions $f$ and $\tilde{f}$ belong to the broader class of all degree-four Morse polynomials with non-degenerate principal homogeneous parts. However, the entire construction proving this proposition in [27] can be performed within the spaces of the versal deformations (1). Thus, it also proves the above Proposition 3.
These constructions provide upper bounds of the numbers of isotopy classes of generic Morse perturbations associated with virtual functions. These bounds are given in Propositions 9, 16, and 21 for classes $X_{9}^{+}$, $X_{9}^{1}$ and $X_{9}^{2}$ separately. Consequently, we obtain estimates of the numbers of connected components of the real discriminant complement which are associated with each virtual component. Specifically, in the cases $X_{9}^{+}$ and $X_{9}^{2}$, all real connected components of discriminant complements associated with the same virtual component can be obtained from each other by the action of a transformation group of order eight. In the case $X_{9}^{1}$ the analogous group is of order four. Then, for each virtual component in the list from [24], we compare all real components associated with this virtual component and determine which of these components are identical. This comparison provides the results of the present work concerning the $X_{9}$ singularities. The strategy for the $J_{10}$ and $P_{8}$ singularities is exactly the same.
@@
In [27], [28], and [29], we use the same strategy with a different definition of virtual components. In these works, we classify the generic perturbations of function singularities up to isotopy within the space of Morse functions instead of the space of non-discriminant functions. Thus, the virtual components of the formal graph are determined there by removing the edges that change the Morse type.
@@
For any generic polynomial $f:{\mathbb{R}}^{n}\to{\mathbb{R}}$, the number $\mbox{Ind}(f)$ equals the number of its real critical points with a negative critical value and an even Morse index minus the number of critical points with a negative critical value and with an odd Morse index.
It is easy to see that this number is an invariant of the connected components of the discriminant complements, and also of associated virtual components.